# Social Good or Scientific Curiosity? Uncovering the Research Framing Behind NLP Artefacts

### v1

@@ 7 Conclusion
In this work, we first proposed a generalizable research framing schema to support consistent interpretation of research intent across NLP domains, then introduced an automated framework for extracting epistemic elements and inferring framings. To assess generalization, we annotated 49 HS papers and found our system outperforms a strong in-context LLM baseline in both AFC and HS. We applied the framework to 102 recent AFC papers to demonstrate its use for scalable literature analysis—revealing a shift toward human-assisted systems, a rise in vague framings, and greater emphasis on justification generation. Future work will explore automated framing and epistemic element discovery in new domains to further streamline NLP research analysis.
@@
  • Generalizability of conclusions – Our findings reveal important trends, such as the prevalence of vague research framings. However, given that our analysis does not cover all NLP papers on automated fact-checking and hate speech detection, these insights should be viewed as well-informed estimates—not definitive claims. Nonetheless, they offer a strong starting point for understanding broader trends in research framing.
  • Tool accuracy – While our model is effective at identifying vague research framings, achieving a precision of 76.16% in fact-checking and 86.24% in hate speech detection, we acknowledge that, as any other automated framework, it still makes mistakes. Individual misclassifications can occur, though large-scale trends remain informative.
  • Appropriate use cases – The tool is designed primarily for scalable analyses of new subfields (where small errors cancel out in aggregate analyses) and writing assistance (offering feedback, research framing rankings, and justification suggestions). It is not intended for peer review or paper acceptance decisions, as it is not precise enough to make definitive judgments about research quality. Furthermore, our aim is not to prescribe how research should be framed, but to surface and clarify existing framings—whether practical, exploratory, or societal—to support reflection on how NLP contributions are positioned and interpreted.
@@ B.1 Semantic clustering via NLI validation
Table 6: Filtered MRR (%) for each epistemic element across sample sizes. DS = Data Subjects, DA = Data Actors, MO = Model Owners, MM = Model Means, AM = Application Means, EN = Epistemic Ends.
@@ C.1.1 Automated fact-checking
[…]
@@
[…]
@@
[…]
@@
Assisted Media Consumption:
$&\big{(}\text{Data subjects: {social media users}}\,\mathbf{\land}\,\text{Model owners: {social media companies}}\\ &\quad\mathbf{\land}\,\mathbf{\bigvee}\big{\{}\text{Application means: {vague persuasion}},\,\text{Ends: {limit misinformation}}\big{\}}\big{)},\\ &\text{Data actors: {media consumers}}\end{aligned}\right.$
Automated Content Moderation:
$&\big{(}\text{Data subjects: {social media users}}\,\mathbf{\land}\,\text{Data actors: {algorithm}}\\ &\quad\mathbf{\land}\,\text{Model owners: {social media companies}}\big{)},\\ &\text{Application means: {automated removal}}\end{aligned}\right.$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ C.1.2 Hate Speech Detection
$&\text{Model means: {humanin the loop}},\\ &\text{Data actors: {social media moderators}}\end{aligned}\right\}$
@@
Automatic Content Moderation:
$&\text{Application means: {supplant human moderators}},\\ &\text{Application means: {automated removal}},\\ &\text{Application means: {generate counter narratives}}\end{aligned}\right\}$
@@
Vague Data Analysis:
$&\text{Application means: {analyze data}},\\ &\text{Model means: {corpora analysis}},\\ &\text{Model means: {data collection}}\end{aligned}\right\}$
[…]
@@
Vague Moderation:
$&\text{Data actors: {social media moderators}},\\ &\text{Data actors: {social media companies}},\\ &\text{Model owners: {social media companies}}\end{aligned}\right\}$
[…]
@@ E.3 Research Framings
Table 7: Ordinal classes for research framings: vague research framings have negative values.

### v2

@@ 7 Conclusion
In this work, we first proposed a generalizable research framing schema to support consistent interpretation of research intent across NLP domains, then introduced an automated framework for extracting epistemic elements and inferring framings. To assess generalization, we annotated 49 HS papers and found our system outperforms a strong in-context LLM baseline in both AFC and HS. We applied the framework to 102 recent AFC papers to demonstrate its use for scalable literature analysis—revealing a shift toward human-assisted systems, a rise in underspecified framings, and greater emphasis on justification generation. Future work will explore automated framing and epistemic element discovery in new domains to further streamline NLP research analysis.
Acknowledgments
Eric Chamoun is supported by an EPSRC-funded studentship. Andreas Vlachos is supported by the ERC grant AVeriTeC (GA 865958) and the DARPA program SciFy. Michael Schlichtkrull is supported by the Engineering and Physical Sciences Research Council [grant number EP/Y009800/1], through funding from Responsible AI UK (KP0016). We’d like to thank Jose Camacho-Collados for feedback on this paper.
@@ Acknowledgments
  • Generalizability of conclusions – Our findings reveal important trends, such as the prevalence of underspecified research framings. However, given that our analysis does not cover all NLP papers on automated fact-checking and hate speech detection, these insights should be viewed as well-informed estimates—not definitive claims. Nonetheless, they offer a strong starting point for understanding broader trends in research framing.
  • Tool accuracy – While our model is effective at identifying underspecified research framings, achieving a precision of 76.16% in fact-checking and 86.24% in hate speech detection, we acknowledge that, as any other automated framework, it still makes mistakes. Individual misclassifications can occur, though large-scale trends remain informative.
  • Appropriate use cases – The tool is designed primarily for scalable analyses of new subfields (where small errors cancel out in aggregate analyses) and writing assistance (offering feedback for authors on the writing, akin to systems such as SWIF2T (Chamoun et al., 2024)). It is not intended for peer review or paper acceptance decisions, as it is not precise enough to make definitive judgments about research quality. Furthermore, our aim is not to prescribe how research should be framed, but to surface and clarify existing framings—whether practical, exploratory, or societal—to support reflection on how NLP contributions are positioned and interpreted.
@@ B.1 Semantic clustering via NLI validation
Table 6: Filtered MRR (%) for each epistemic element across sample sizes. DS = Data Subjects, DA = Data Actors, MO = Model Owners, $M$ = Model Means, AM = Application Means, EN = Epistemic Ends.
@@ C.1.1 Automated fact-checking
$\land\,\text{Model means: {generate claims}}$
$\land\,\text{Application means: {produce misinformation}}$
@@
$\mathbf{\bigvee}\big\{\text{Data subjects: All except {professional journalists}, {citizen journalists}, {technical writers}}\big\}$
$\land\,\mathbf{\bigvee}\big\{\text{Data actors: {professional journalists}},\textit{citizen journalists},\textit{technical writers}\big\}$
$\land\,\mathbf{\bigvee}\big\{\text{Application means: All except {supplant human fact-checkers}, {not specified}}\big\}$
$\land\,\mathbf{\bigvee}\big\{\text{Ends: {limit misinformation}},\textit{limit AI-generated misinformation}\big\}$
@@
$\mathbf{\bigvee}\big\{\text{Data subjects: {professional journalists}},\textit{citizen journalists},\textit{technical writers}\big\}$
$\land\,\mathbf{\bigvee}\big\{\text{Data actors: {professional journalists}},\textit{citizen journalists},\textit{technical writers}\big\}$
$\land\,\mathbf{\bigvee}\big\{\text{Application means: All except {supplant human fact-checkers}, {not specified}}\big\}$
$\land\,\mathbf{\bigvee}\big\{\text{Ends: {increase veracity of published content}},\textit{limit misinformation}\big\}$
@@
$\mathbf{\land}\,\mathbf{\bigvee}\big\{\text{Model means: {classify/score veracity}},\textit{evidence retrieval},\textit{human in the loop}\big\}$
$\mathbf{\land}\,\text{Ends: {increase veracity of published content}}$
Assisted Media Consumption:
@@
$\Big(\text{Data actors: {scientists}}\land\text{Model owners: {scientists}}\Big)$
$\lor\,\text{Ends: {develop knowledge of NLP/language}}$
@@
$\land\,\mathbf{\bigvee}\big\{\text{Model means: {evidence retrieval}, {classify/score veracity}}\big\}$
$\land\,\mathbf{\bigvee}\big\{\text{Application means: {gather and present evidence}},\textit{provide labels/veracity scores}\big\}$
@@
$\land\,\text{Application means: {identify claims}}$
@@
$\land\,\mathbf{\bigvee}\big\{\text{Ends: {limit misinformation}},\textit{limit AI-generated misinformation}\big\}$
@@ C.1.2 Hate Speech Detection
[…]
@@
$\land\,\text{Data actors: {engineers and curators}}$
Automatic Content Moderation:
@@
$\land\,\text{Ends: {detect hate/falsehood for law enforcement}}$
Vague Data Analysis:
@@
$\land\,\text{Ends: {fight hate}}$
@@
$\land\,\text{Application means: {identify toxic content}}$
Vague Moderation:
@@
$\land\,\text{Data subjects: {social media users}}$
$\land\,\mathbf{\bigvee}\left\{\text{Application means: All except {supplant human moderators}, {automated removal}}\right\}$
$\land\,\mathbf{\bigvee}\left\{\text{Model means: All except {human in the loop}}\right\}$
@@ E.3 Research Framings
Table 7: Ordinal classes for research framings: underspecified research framings have negative values.